Michael Faraday unsuccessfully attempted to measure the potential difference across the River Thames due to the flow of water (which is a good conductor due to dissolved salts) in the earth's field. At Waterloo Bridge, where Faraday made his attempt, the river is 200 m wide and flows eastward at 2.4m/s . The earth's magnetic field here has a magnitude of 50 ?T and is tipped 65?from the horizontal, so it has both horizontal (north) and vertical (down) components.

Part B

What is the magnitude of the potential difference?
Answer #1

The magnitude of the potential difference is,

\(\varepsilon=B l v \sin \theta=\left(5 \times 10^{-5} \mathrm{~T}\right)(200 \mathrm{~m})(2.4 \mathrm{~m} / \mathrm{s}) \sin 65^{\circ}=2.2 \times 10^{-2} \mathrm{~V}\)